OpenStudy (anonymous):

someone please help!! In a small bag of M&Ms, there are 9 orange candies and 4 blue candies. You select 1 candy, eat it, and then select another. Find P(you select first an orange candy and then a blue candy). Round your answer to the nearest hundredth.

OpenStudy (anonymous):

wel P( you get orange first)=9/13 P(you get blue after)=4/12( the 12 as you loose on m&m you ate it!) so now 9/13 * 4/12= something, get a calculator

OpenStudy (anonymous):

Thankyou!